OpenStudy (anonymous):

A student used a balance and a measuring cylinder containing water to find the density of a small stone. mass of stone = 20g original volume of water in cylinder = 30 cm^3 volume of stone and water = 38 cm^3 Density of stone = ________

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the volume of the stone, is: 38-30=...cm^3

OpenStudy (anonymous):

8 cm^3

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

so, the requested density, is: \[\Large \delta = \frac{m}{V} = \frac{{20}}{8} = ...\frac{{{\text{grams}}}}{{{\text{c}}{{\text{m}}^{\text{3}}}}}\]

OpenStudy (anonymous):

2.5

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

thats the ans?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

yes! density= 2.5 grams/cm^3